OpenStudy (anonymous):

FLVS problem and i have no clue how to finish this problem Solve 4x2 = 12x + 40.

OpenStudy (anonymous):

4x2 = 12x + 40 4x^2 - 12x - 40 = 0 Get rid of 4 by taking commons 4(x^2 -3x - 10) = 0 Divide 4 on both sides (x^2 -3x - 10) = 0 Now factor: x^2 - 3x -10 =0 x^2 -5x + 2x - 10 = 0 take commons: x(x-5) +2(x-5) = 0 (x+2)(x-5) = 0 solve them for Zero: x = -2 x = 5

OpenStudy (anonymous):

taking 4 common on both sides.. x^2 = 3x + 10 x^2 - 3x - 10 = 0 now factorize x^2 -5x + 2x - 10 = 0 x(x-5) +2(x-5) = 0 (x-5)(x+2)=0 Either (x-5)=0 this gives x=5 Or (x+2)=0 this gives x= -2

OpenStudy (anonymous):

@iveymarie got it..??

OpenStudy (anonymous):

Thankyou @Kashan and @Skysthelimit I was getting so stuck at the factoring part!(:

OpenStudy (anonymous):

@iveymarie your welcome!! I'm in flvs to